OpenStudy (anonymous):

how would you factorise x^4 -x = 0 x(x^3 -1)=0 x(x-1)(........?.)=0 i cant seem to go further than that, wouldnt the blank be x^2 +1

OpenStudy (mathmale):

Thank you for showing what you've done so far. You're exceptional! x(x^3 -1)=0 x(x-1)(........?.)=0 are both fine. x^3-1 = x^3 - 1^3 is "the difference of two cubes." Where have you seen that name before? Hint: Look up "special products and factors."

OpenStudy (anonymous):

my textbook says x(x-1)(x^2 +x +1) =0 i dont get how you get (x^2 + x +1)